Assume that the variable is normally distributed, the average time it takes a group of senior high school sidents to complete a certain examination is 40.2 minutes while the standard deviation is 8 minutes


What is the probability that a randomly selected senior high school students will complete the examination in less than 43 minutes? the given altuation, what words would give us a clue to the direction of the


curve or where the shaded part is located?


1
Expert's answer
2022-06-13T09:53:07-0400
"P(X<43)=P(Z<\dfrac{43-40.2}{8})"




"=P(Z<0.35)=0.6368"